OpenStudy (anonymous):

Israel started to solve a radical equation in this way:

OpenStudy (anonymous):

what mistake did isreal make?

OpenStudy (anonymous):

Israel started to solve a radical equation in this way: Square root of x plus 6 − 4 = x Square root of x plus 6 − 4 + 4 = x + 4 Square root of x plus 6 = x + 4 (Square root of x plus 6)2 = (x + 4)2 x + 6 = x2 + 8x + 16 x + 6 − 6 = x2 + 8x + 16 − 6 x = x2 + 8x + 10 x − x = x2 + 8x + 10 − x 0 = x2 + 7x + 10 0 = (x + 2)(x + 5) x + 2 = 0 x + 5 = 0 x + 2 − 2 = 0 − 2 x + 5 − 5 = 0 − 5 x = −2 x = −5 Solutions = −2, -5

OpenStudy (anonymous):

A) He subtracted 6 before subtracting x. B) He added 4 before squaring both sides. C) He factored x2 + 7x + 10 incorrectly. D) He did not check for extraneous solutions.

OpenStudy (anonymous):

@abb0t can you help?

OpenStudy (abb0t):

Can you write the function using the LaTex equation editor on the bottom of this?

OpenStudy (abb0t):

Which do you think it is.

OpenStudy (anonymous):

√x+6 -4=x

OpenStudy (anonymous):

im not sure im confused.. could you help?

OpenStudy (anonymous):

@phi caan you helpp?

OpenStudy (anonymous):

@whpalmer4 help? :)

OpenStudy (anonymous):

@dmezzullo help :)

OpenStudy (anonymous):

@timo86m help:)

OpenStudy (anonymous):

it be better if you wrote it or used perentesis LIke sqrt(x)-4 is different from sqrt(x-4)

OpenStudy (anonymous):

√(x+6) -4=x

OpenStudy (anonymous):

better? @timo86m

OpenStudy (anonymous):

help.. @timo86m

OpenStudy (anonymous):

C) He factored x2 + 7x + 10 incorrectly. I wanna say that check it yourself :) are the factors of it (x + 2)(x + 5) easiesst way is to expand (x + 2)(x + 5)